OpenStudy (anonymous):

find the solution set: \[\sqrt{2x - 6}\]=\[\sqrt{5x }\] - 3

OpenStudy (anonymous):

find the solution set: \[\sqrt{2x-6}\]=\[\sqrt{5x}-3\]

OpenStudy (anonymous):

have you tried squaring both sides?

OpenStudy (anonymous):

yes. i'm trying it right now :)